OpenStudy (anonymous):

How would you solve this 2x – 3y = –29 x + 4y = 13 using substitution?

OpenStudy (anonymous):

2x-3y=-29 x+4y=13 4x-6y=-58 4x+16y=52 -22y=-110 y=5 x+20=13 x=-7

OpenStudy (anonymous):

Would this 2(4y + 13) – 3y = –29 be the new equation after the expression equivalent to x from the second equation is substituted into the first equation?

OpenStudy (anonymous):

idk substitution is for pansies

OpenStudy (anonymous):

Lol... But I need to know.